LEVEE   PROGRESS   REPORTS   (LPR'S) 

(revised/updated 11/5/05)

 

MOPR Notes:   All are submitted by DC, Lev Project Coordr/Consultant unless otherwise noted.  All "Subjects" begin with "Progress Coordination Report on" but are referred to as LPR's.

LPR's posted below with the oldest atop are those dated: 

 


 

To:  VP Levee Cmsn  -  12/21/98  

Subj:  Levee & Structures Item 3A, VP, MO

Progress as of 11/25/98 - NorVel continues to disk & aerate the on site fill material & place impervious fill with scrapers in portions of area 10 & 11. Trucking fill to area 9 between sta 107+00 & 111+00.  

SCI {sic} has been on site with a technician monitoring impervious fill placemt & testing for moisture & compaction.  NorVel has completed inspection trench & impervious fill prep work on south side of St Louis Ave Closure Structure. 

 

A hazardous waste contractor ((ARDL?)) visited the job site to inspect the barrels that have been found & secured on site.  No tests of contents were performed at this time.  Contractor visually inspected the location & condition of the barrels in preparation for another visit at which time the contents of the barrels will be tested.

 

Progress as of 12/2/98 - NorVel continues placing impervious fill material between sta 114 to 119 & 123 to 131.  Trucks are hauling material from Simpson Quarry Burrow ((offsite)) Site & placing impervious fill between sta 106 & 110.  Contractor worked on Fri & Sat after Thanksgiving & placed over 6300 cy of fill on the 2 days combined.

 

SCI has been on site during these ops to monitor the fill placemt & make moisture/compaction.  The contractor has removed an add'l 177 tires from the site & disposed of off site.  A total of 374 tires have been removed from site & disposed of.  NorVel has started placing fill at sta 122 + 50 to an elevation of about 118.  This allows him access to a longer stretch of levee to transport, place & aerate impervious fill scraper dirt.  Hazardous waste contractor ARDL, Inc were on the site 12/1/98 under direction of COE & obtained samples of domestic trash from inspection trench excavation near sta 131 + 25, Lagoon sludge & 5 barrels of unknown liquid materials that contractor is storing on site.

 

Progress as of 12/9/98 - This week, NorVel continued placing impervious fill material by both truck & scraper in areas 10 & 11.  Placed trucked dirt in areas 8 & 9 south of St Louis Ave gate structure.  No safety violations or injuries reported this week.  No results from ARDL, Inc testing that took place last wk.

Progress as of 12/16/98 - This wk, NorVel continued placing impervious fill material by both truck (offsite burrow) & scraper (onsite borrow).  The fill is being placed in areas 8, 9, 10 & 11 which relates to sta 104 + 50 to 131 + 90 skipping areas between sta 111 + 50 to 113 + 50 & everything past sta 131 + 90 due to pending mod work in those areas.  The largest portion of fill is being placed between sta 104 + 50 & 122 + 50.  Approximate total of impervious fill placed this wk was 11,500 cy. >  DC.

 

_____________________________________________________________

To:  VP Levee Cmsn - 1/19/99  

Subj:  Levee & Structures Item 3A, VP, MO

As of 12/23/98 - On 12/17 & 18, NorVel coninued placing impervious fill material in areas 8, 9, 10 & 11.  On 12/21 thru 12/23, all placing of fill was halted due to rain, freezing rain & snow.  Mosley Const is still awaiting COE to submit P&S for the mod on the Simpson Lake Gravity Drain.  No results from ARDL, Inc on testing that took place on 12/1/98.  No safety violations or injuries reported this wk.  The COE cancelled coordn mtg for 12/23/98.

As of 12/30/98 - NorVel was unable to place any impervious fill this week due to weather conditions.  Mosley Const continues to revise schedules which reflects the effects the pending mods will have on the progress of Item 3A.  No results from ARDL, Inc on testing that took place on 12/1/98.  COE cancelled coordn mtg on 12/30/98.   >  DC.

_____________________________________________________________

To:  Mayor, City Atty & BOA - 2/1/99

Subj:  Levee & Structures Item 3A, VP, MO

As of 1/6/99 - All controlling work is on hold pending mod negotiations or has been stopped due to weather conditions.  Mosley Const has submitted updated progress to COE for their review.  Still waiting for P&S for mods to the Simpson Lake Gravity Drain.  No results from ARDL, Inc on testing that took place on 12/1/98. 

As of 1/13/99 -  The coordn mtg was cancelled due to adverse weather conditions.

 

As of 1/20/99 -  All controlling work is still on hold due to adverse weather conditions.   Mosley Const continues to revise schedules which reflects the effects the pending mods will have on progress of Item 3A.  Results from ARDL testing of waste material we rcv'd & are to be handled as special waste.  The material will be hauled to "Superior Landfill, Inc" for disposal.  New applications for disposal of special waste from Item 3A will have to be filled out & signed by the Mayor.  Test results for contents in the 5 barrels were not rcv'd.  COE is still working on design of the gravity drain into Simpson Lake. 

 

As of 1/27/99 -  All controlling work is still on hold due to wet soil conditions.  Contractor began the care of water op by starting to pump water out of Simpson Lake Det Area.  COE is completing design of box culvert & outlet structure for Simpson Lake Gravity Drain. 

Levee coordr has contracted {sic} several hazardous waste contractors, requesting cost proposals for their svcs to test, remove & dispose of any non-hazardous, special & hazardous wastes encountered during const of the levee.

 

On 1/25/99, Lev Coordr, Lev Op & Maint Team, PBW Dir, Emerg Mgmt Coord, Bldg Cmsnr & St L Cnty POL Dept supervisor met with COE staff & reviewed the draft copy of op procedures for Trolley Gate at Vance Rd closure structure.  A hands-on training op of gate is being planned for end of February. >DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 5/15/00

Subj:  Levee & Structures Item 2 & 3B

As of 4/28/00:  (Item 3B) - On 4/15 I conducted a group of City ofcl's & Lev Cmsn members on a tour of g/p ruins.  Purpose was to show them the extent of these old foundations & underground chambers which will have to be removed, demolished & hauled to a land fill in order to complete final item of levee.

On 4/17 KAJACS were given Notice To Proceed with const of Item 3B.  They are clearing & rocking the pkg lot on NW corner of 6th & Glover in order to setup their op trailers.  They will begin to mobilize their equipmt on this site later this wk.  They have also started clearing the area along const limits in order to erect safety fencing.

 

On 4/24 KAJACS completed erecting safety fencing along const limits.  They have setup their const ofc trailers (2) at 6th & Glover & erected a temp chain link fence around this site.  They have started a tire & battery collection area on this site.  Contractor started clearing certain areas so they can begin excavating for their inspection trenches.  On 4/26 Norvell began to mobilize their equipmt & start clearing, grubbing & stripping at West const limits.

 

As of 5/5/00:  KAJACS has completed the setting up of their const offices.  They continue to excavate, inspect & backfill inspection trenches.  They also continue collecting tires (80) & car batteries (1) from the site.  Norvell continues clearing from the West limit to the East.  They have also rcv'd approval of locations for their burn pits.

 

As of 5/12/00:  (Item 3B & 2) - KAJACS continues to excavate inspection trenches.  Inspection trench from sta 163 + 50 to sta 155+00 has been excavated.  Approx 1200 LF {sic} of inspection has {sic} excavated.  Due to 3.60" of rain on 5/7/00 most excavating has been delayed except for drainage ditches.

Norvell have cleared approx 5 acres of trees on North side of creek & approx 15 acres of trees on South side of creek.  Clearing op has been delayed this wk due to rain & high water.  Norvel has excavated their burn pit & placed their burning equipmt in the pit.  They will start the burning op as soon as trees dry sufficiently to burn.

An onsite mtg at Fishpot Lev (Item 2) was held on 5/11/00 @ 9am with reps from COE & myself to solve crown vetch & erosion problems on slopes of South end of levee.  The results of this mtg will be reported at a later date.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 6/17/00

Subj:  Levee & Structures Item 3B

As of 5/19/00:   KAJACS Inc continue draining, drying out & then compacting inspection trenches from sta 164 to 153;  & continue removing debris from the site, having collected 96 tires to-date.  Norvell has started the buring op & have burned approx 10 acres of dried clearing material from South side of GG Creek. 

 

As of 5/26/00:  KAJACS continues drying out inpsection trenches & then compacting them from sta 164 to 153 & have started excavating & backfilling inspection trenches to East from sta 153 to 147 & excavating inspection trench from 145 to 139.  They also have removed all asphalt road material within const limits.  They continue removing tires & other debris & stockpiling it on site.  NORVEL continues their clearing & burning op.  They have completed burning approx 15 acres on South side of creek. 

 

As of 6/2/00:  KAJACS continues drying out & then compacting inspection trench from sta 165 to 148.  They continue excavating & backfilling inspection trench from sta 147 to 143.  They continue removing debris & stockpiling on site.  NORVEL continues clearing & burning ops.  They have burned approx 20 acres on South side of creek & cleared approx 10 acres on the North side of it.

 

KAJACS are removing concrete rubble from the East end (behind the LeLu Bldg) of site. This rubble will be stockpiled & disposed of on site.  They have also discovered black top rubble, which will have to be handled as special waste & hauled off site & disposed of.  KAJACS is excavating some inspection trenches to determine the extent of this type rubble.  They have also started bldg the lev from sta 154 +00 to sta 148 + 00.  They are excavating the storm water det basin.  They are stripping & stockpiling top soil on South side of creek.  They have started to excavate the new creek channel from sta 154+00 to sta 149 + 00.

 

NORVEL has completed grubbing, clearing & burning op on South side of creek.  They continue grubbing & clearing on North side of creek.  They have moved the burn pits & are now burning on North side of creek.  NORVEL has started bldg the lev from sta 166 + 50 to sta 154 + 00.  They are also excavating the storm water det basin.   >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 5/21/01

Subj:  Levee & Structures Item 3B

 

As of 4/13/01:  Kajacs Contractors, Inc have started placing aggregate on levee road & access ramps.  They continue to haul, place & stockpile bedding aggregate 400 # riprap.  Subcontractor has completed fertilizing, seeding & mulching land side of levee slopes, det pond, mit area & earth {sic}.  Stockpile at Simpson Lake det pond.

As of 4/20/01:  Kajacs hauling, placing & stockpiling of bedding aggregate & 400 # riprap was halted this wk due to wet soil conditions.

As of 4/27/01:  Kajacs cont'd hauling, stockpiling & placing bedding aggregate & 900 # riprap this wk. 

As of 5/4/01:  Kajacs have completed hauling & stockpiling bedding aggregate & riprap materials.  They continue placing bedding & riprap materials & have started finish grading slopes on creek side of levee above riprap for subcontractors to start seeding op.

As of 5/11/01:  Kajacs continue placing bedding & riprap on creek side of levee; started to close haul road gap in levee & continue to finish grade levee slopes.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 5/20/02  

Subj:  Levee & Structures

On 4/18/02, an on site inspection was made by Gary Camp, John Boeckmann of USACOE Hydraulic Branch Regulatory Section & myself.  This inspection was done to familiarize Mr. Boeckmann who will be replacing Mr. Camp who is retiring in 10/02.  Briefed Mr. Boeckmann on our emerg op plan using certain sections of lev during a flood.

On 4/23/02, I met with Jonathan Baer, an env'l scientist for DG Purdy & Assoc, Inc.  We did a very thorough tour of completed sections of lev & of ROW sections of Item 4B.  Mr. Baer is working on a plan & design for env'l mitigation on the project. Submitted by DC, Lev Project Coordr/Consultant.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA  -  9/16/02  

Subj:  Glass Plant Clearing Contract

On 8/14/02, an on site inspection of g/p Ruins was made by the COE, Boone Const, StLCnty DOH/Waste, PH Weis & Assoc & City VP to work up a purchase order contract to remove all standing timber (excluding marked bat trees), fallen timber debris & stockpile on site.  All special waste & man-made objects shall be stockpiled on site in separate locations.  The above site contains approx 10 acres. 

On 8/16/02, a purchase order contract was awarded to Boone Const by the StLDist COE Contracting Ofcr.  On 8/20/02, Boone Const & their sub-contractor, Magurder Excavating, began to mobilize equipmt & start clearing the g/p site.  On 8/24/02, Magurder Excavating had completed their part of clearing & stockpiling.  Boone is still working on clean-up special waste & man-made objects at the site.> DC

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA  -  3/17/03  

Subj:  Item 4B (Phase 2)

Tree clearing contractor, Ceres Env'l Svcs, Inc, started clearing trees east of BNRR tracks on 2/24/03.  Tree clearing is proceeding to east, weather permitting.  Contractor is in the process of obtaining their burn permit from DNR.  On 3/7/03 a mtg was held at VP with SBC.  This mtg was coord'd by COE with SBC & city ofc'ls.  We are trying to determine the location of the 54 "0 {sic} storm water pipe from Old Mer Sta Rd to Fishot Det Pond.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 7/21/03  

Subj:  Levee & Structures.

Wk of 7/2/03, a pre-bid & site visit conf for levee & structures on Item 4B was held at city hall.  Numberous contractors & myself were in attendance (see attached attendance roster).  This conf was conducted by COE.  9 site locations were visited (see attached list).   Wk of 7/7/03:  On 7/10/03 the contractors resumed burning material.  Tire removal should begin on 7/11/03.  >  DC  >  MOPR Note:  The above-mentioned roster & list were not in the BOA pkt that contained this LPR.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 11/17/03  

Subj:  Item 4B (Phase 3)

Wk of 10/13/03:  At a 10/16/03 pre-const mtg with Contractor ESI, COE & City VP, various phase of P&S were discussed.  Also at a 10/16/03 mtg with numerous utility reps, COE, ESI & CIty VP (see attached sign-in sheets for each mtg) ((below)), schedules for relocation & removal of utilities were discussed & planned.  Wk of 10/20/03:  American Electric shut the electric supply to sport complex off.

Wk of 11/3/03:  VP PBW personnel started removing chain link & wooden fences inside levee footprint of the sport complex.  Wk of 11/10/03:  JM & I will be working on bid specs for asbestos removal & demolition of structures in the AL area.  >  DC:rp

 

2 Sign-In Sheets:  ((Ref PCO List for phone #'s both attachmts))  Preconst Conf VP Levee & Structures Item 4B, VP, MO Contract DACW43-03-C-0418, ESI:  ((10/16/03))  Mike Feldmann, Vick James, Len Ross, all of COE Const;  Alan E. Wolfe, Danny Thompson, Jim Munchrath all of ESI;  JZ, COE PM ((Project Mgr))Pat Conroy, COE ED ((presume Eng Design))Greg Bertoglio, COE;  DC, City VP;  CLM, McKinney Assoc  - VP.

10/16/03 Utility Relocation Mtg:  ((All same as above +)) TW, EM, Craig Donis, COE R.E.;  Carol Kreutzer, COE, CE ((for Corps Eng or ?))Rick Olson & Lisa Wright, MAWC; Gary Drikow, Lac Gas;  Christine Grunbaum & Shelley Morris, AmerenUE;  Sherry Sch(mi?)dle, SBC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA  -  12/15/03 

Subj:  Item 4B (Phase 3)

Wk of 11/17/03:  VP PBW personnel completed the removal of chain link & wood fences, dugouts, etc.  The only remaining items to be removed are the light standards & some misc items at the Sports complex.   Wk of 11/24/03:  VP PBW dir has set up a mtg with an interested party to remove the light standards at the Sports Complex. The  mtg will be on 12/1/03 at 9:30, JM, myself & interested party will attend.  Wk of 12/1/03:  Contractor ESI was given the Notice to Proceed & has started to mobilize his equipmt at the Sports Complex weather permitting & will probably start to layout const limits & control points.  On 12/5/03, ESI, Len Martin COE, Pat Palmer Ops Div Mgr STL Cnty Hwy & Traffic, JM & myself video taped the condition of the streets that will be used as haul roads prior to levee const. >  DC:rp

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA  -  2/17/04 

Subj:  Item 4B (Phase 3)

Wk of 12/29/03:  Contractor has located the const ofc trailers @ 101 Marshall Rd (at a site on Valley Material prop) & has erected temp project safety fencing as required by the safety & health requiremts manual.  Wk of 1/5/04:  Contractor is excavating & grading an area at the southeast corner of the g/p ruins for the staging area & the crusher location.  Wk of 1/12/04:  Contractor has started removing the wash out concrete debris & hauling larger pieces of concrete to the staging area.  City has acquired all the props in AL & has prepared bid specs for the demolition of all the houses in AL.  Bid specs were advertised with bid openings on 1/30/04.  Subcontractor started clearing in various areas of project.  Ameren UE has cont'd relocating the utility lines along River Dr.  The wkly progress mtg with ESI, COE & city reps started on 1/15/04.  

 

Wk of 1/19/04:  ESI continues to remove the wash out concrete debris & haul to crusher staging area.  Subcontractor continues clearing in Det Area B, 5th St gravity area & riverside borrow area east on Pharoah.  ESI has laid out the toe of levee slopes in Sportsplex.  Subcontractor (Loana Contractors, Inc) will set up a const ofc trailer on the Lions Club site.  Wk of 1/19/04:  City VP PBW crew removed the light standards within the levee footprint in the Sportsplex.  Construction ofc trailers are now completely op'l.  Wk of 1/26/04:  ESI continues to remove the wash out concrete debris & haul to the crusher staging area & have also started to remove g/p concrete foundation walls & brick structures.  Subcontractor continues clearing & grubbing ops as weather permits. 

 

On 1/28/04, Mississippi Valley Div Cmdr Brigadier Gen Don T. Riley visited VP to see the levee project 1st hand.  At 3:30pm the General, COE personnel & VP ofc'ls met at city hall & toured the flood prone area & completed levee segmts within City VP. COE personnel that accompanied the Gen were Major Joseph Tyron, JZ, Pat Conroy, Julie Ziino & Greg Bertoglio.  City ofc'ls present were Mayor DM, Jim Mitas, EM, CLM, Laura Kinder (Asst Supt of VP Schools), Jim Probert (VP School Bd & Levee Cmsn Member), DA (Ald 1st Ward) & myself.  At 4:15pm the Gen, COE personnel & city ofc'ls met with COE Const personnel Mike Feldman {sic}, Vick James, Len Martin & Alan Wolfe, Gen Mgr of ESI, at the const site to observe work in progress.  At 4:45 pm the Gen, COE personnel & city ofc'ls returned to city hall to view flood photos & discuss issues.  The purpose of this project visit was to provide Gen Riley a 1st hand understanding of VP flooding situations, levee progress & the need for add'l fed funds in FY 04. 

 

On 1/29/04, COE Const personnel, ESI & city ofc'ls met with MAWC to coord locations, relocations & abandonmts of water lines within the footprints of levee.  Various water line designs were discussed & on site mtgs were scheduled with MAWC, contractor & city ofc'ls.  At 3pm on 1/30/04 at city hall, the demolition bids for vacant structures at AL were opened.  Bid amts & Award of Bids will be announced at a later date.  Wk of 2/2/04:  On 2/2, the bid for the demolition of the structures at AL was awarded to Aherns Co for $29,439.00.  This contract will be completed by 3/1 (see attached bid list).  On 2/3 a mtg was held at city hall with the affected bizs in BPlz & concerns by biz owners were answered by the contractors.  (see attached info letter, plan profile drawing & attendance sheet) 

 

On 2/6/04, a mtg was held at the const ofc trailer with AmerenUE & COE personnel, ESI & myself (see attached roster).  Completed relocations were reviewed.  Schedules for upcoming relocations to coord with contractor schedule for levee const.  All AmerenUE work has been completed east of BNSF RR tracks.  This mtg has ESI & AmerenUE working on the same priorities to complete various stages of const.  ESI has mobilized their crusher on site & weather permitting they have started crushing wash out concrete debris & g/p foundation debris.  Clearing & grubbing continue weather permitting.  Work has slowed some due to weather conditions.  Done {sic} time allows many loose ends to be brought together.  DC:rp  Enclosures.  ((will try to get & post bid list, letter, drawing & 2 rosters))

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA  -  3/15/04 

Subj:  Item 4B (Phase 3)

 

Wk of 2/9/04: Contractor ESI continues to remove washout concrete debris & g/p foundation & brick debris.  Clearing & grubbing continue at 3rd & 5th Streets gravity drain locations.  Work is progressing very rapidly even with nasty weather conditions.  

 

The wk of 2/16/04:   ESI has completed removal of the major portion of washout concrete debris.  Minor cleanup remains.  They continue to remove & crush g/p foundation & brick debris & stockpile & to stockpile special waste (tires, metal & clearing debris) in Sportplex.  Clearing contractor is clearing & grubbing the area East of Pharoah & South of Pyramid Properties, & has installed project safety fencing along const limits south of Pyramid.  Snow & ice have caused some delays this wk.       >  DC:rp

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 4/19/04

Subj:  Item 4B (Phase 3)

Wk of 2/23/04:  ESI continues:  to excavate & crush g/p foundations & brick structures & stockpile;  to stock pile special waste materials in the Sportsplex clearing;  clearing & grubbing.  City VP PBW Dept is working to locate the sanitary sewer of Fly By Liquor Store in BPlz;  also with City Eng TW to locate & relocate sanitary sewer lines at bizs located west of Mer Sta Rd adjacent to AL.  Env'l Consultants, LLC is inspecting AL structures for hazardous materials.

 

Wk of 3/1/04:  ESI continues to work in g/p east of BNSF RR ROW in Sportsplex area & east of Pharoah.  Weather delays 3/3, 4 & 5.  City PBW JM, Cmty Dev Dir JEM, City Eng TW & Bom Const Co temp'ly relocated sanitary sewer lines for bizs west of Mer Sta Rd adjacent to AL. JM, MSD & City Hired Inspection Co located sanitary sewer lines of Fly By Liquor Store will not interfere with 54"0 {sic} RCP storm water line thru BPlz.  AmerenUE removed electric lines & poles. 

Wk of 3/8/04:  ESI continues work in g/p, east of BNSF RR ROW in Sportsplex & at 3rd & 5th Sts outlet structures.  This work was delayed by high water 3/8 thru 12.  Midwest Asbestos removed hazardous materials from AL strucutres 3/8 thru 3/12.  Demolition Contractor Aherns Co began work in AL 3/11 & completed it 3/13.  MSD removed the lift sta controls, pumps & metal structure at ground level.  All gas, phone, water & sanitary svcs have been abandoned. 

 

Wk of 3/15/04:  ESI continues work from east of BNSF RR ROW to west of Pharoah Ave & at the 5th St outlet structure.  Bom Const has started const on new lift sta at Marshall Rd & Kena St.  MAWC grouted the water main in AL.  All work required by city in AL has been completed without any delay to contractor.  (Good job by all!)

 

Wk of 3/22/04:  ESI continues work from east of BNSF RR ROW to west of Pharoah Ave & the 5th St outlet strucutre.  Bom continues work on the lift sta at Marshall Rd & Kena.  City continues to work with MerPlz owner on relocation of water line in alley, south of the plaza. 

 

Wk of 3/29/04:  ESI is working on the Levee from Sta 71 +00 to  Sta 87.50 (inside the Sportsplex).  Work has been delayed due to weather from 3/29 thru 4/1.  City continues trying to work with MerPlz owner (without much success) on water line relocations.  If this is not resolved soon it will start a delay in const.  Wk of 4/5/04:  ESI continues to work on 3rd & 5th Sts outlet structures & has removed all the concrete foundations, slabs & asphalt from the st & driveways in the AL area;  also completed add'l clearing & grubbing in AL.  They have pumped & drained the area south of the homes on Pyramid east of Pharoah & west of Kena.  They continue to work in the area west of Pharoah & east of BNSF RR ROW.  City PBW Dept completed their work in AL area with the filling of abandoned life {sic} sta with concrete.  City continues working with MerPlz owner on water line relocations (some progress is being made).  Bom  has started to install the perm sanitary lift sta for the biz located west of Mer Sta Rd & south of Marshall Rd. >  DC:rp

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 6/21/04

Subj:  Item 4B (Phase 3)

 

Wk of 5/10/04:  ESI continues to work on the levee from Sta 72:00 (west end of the Sportsplex) to Sta 98:00 (Kena & Pyramid) & with g/p excavation & blending of eng'd fill.  The work on 3rd & 5th St storm water outlet structures is delayed by high water levels & rain.  City continues to coord with the utility companies to abandon & relocate their facilities within the levee const limits & to work with Mer Plz owner on utility relocations.

 

Wk of 5/17/04:  ESI continues to work on the levee from Sta 72:00 (west end of the Sportsplex) to Sta 98:00 (Kena & Pyramid) & with g/p excavation & blending of eng'd fill as weather permits.  Loana continues to work on the 5th & 3rd St outlet structures & are also placing the 54" ((0 is crossed out)) RCP across River Dr for the 5th St storm water drain to outlet structure as weather & river level permits.  City continues to coord with utility companies to abandon & relocate their factilities within the levee const limits, & also work with MerPlz owner on utility relocations.  City PBW Dept is working to locate city owned abandoned sanitary sewer & water lines.  To date there have been 9 weather delay days for May. 

 

Wk of 5/24/04:  On Mon, 5/24/04, ESI & Loana began video taping the interior of the biz & exterior of the bldgs in BPlz that will be affected by const of the 54" ((0 is crossed out)) storm water pipe thru the plaza.  City PBW Dept has located & prepared city owned sewer & water lines for grouting & will complete as weather permits.  Work has begun on relocation of water lines into MerPlz.  Ameren UE will begin removing old poles, setting new poles & relocating electric svc to MerPlz as weather permits.  Const for May has been very slow with only 4 complete working days in May.

 

Wk of 5/31/04:  ESI continues to work on the levee from Sta 72:00 to Sta 98:00 with g/p excavation & blending eng'd fill as weather permits.  Subcontractor Loana continues to work on the 5th & 3rd St outlet structures as weather & river level permits.  Contractor Kokash {sic} Plumbing is relocating the water line into MerPlz.

 

Wk of 6/7/04:  ESI continues to work on the levee from Sta 72:00 to Sta 98:00 with the g/p excavation & blending eng'd fill material.  Sub-Loana continues to work on the 5th & 3rd St outlet structures as weather & river levels permit & has started working in BPlz.  Kokash Plumbing continues to relocate water line into MerPlz.  MAWC is placing water tap under Marshall Rd to complete the relocation of the water into MerPlz.  AmerenUE started relocating the poles behind MerPlz & relocating the power lines into it. >DC.  

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 9/21/04

Subj:  Item 4B (Phase 3)

 

Wk of 8/9/04:  ESI continues with waste excavation, crushing & blending eng'd fill material.  They continue to work the lev from sta 40+00 to sta 52+00;  from sta 72+00 to sta 98+00;  from sta 34+60 to sta 38+00;  & in the AL Area sta 26+00 to sta 32+00.  They also continue to haul levee material from Simpson Lake Det area & from Simpson Material.  Loana will continue working on headwall for Outlet Structure into Fishpot Det area. 

 

Wk of 8/16/04:  ESI continues with concrete excavation, crushing & blending eng'd fill material.  They continue to work the levee fill from sta 26+00 to sta 32+00;  from sta 40+00 to sta 52+00;  & from sta 67+00 to sta 98+00.  They also continue to haul lev material from Simpson Lake Det area & from Simpson Material.  Loana & Jacobsmeyer continue to work on headwall for Outlet Structure into Fishpot Det area & at Junction Box 1 (between Mer Sta Rd Hwy 141 at BPlz.)

 

Wk of 8/23/04:  ESI continues with waste concrete excavation, crushing & blending eng'd fill material.  They continue to work the levee fill from sta 26+00 to sta 32+00;  from sta 40+00 to sta 52+00;  & from sta 67+00 to sta 98+00.  They also continue to haul levee fill material from Simpson Lake Det area & from Simpson Material.  They have dug inspection trenches from sta 98+00 to sta 99+20 & from sta 99+60 to sta 102+00.  Loana & Jacobsmeyer continue to work 54" gravity drain at BPlz.  Sub-contractor Kelpe started rip-rap, bedding & C-stone at the 141 gravity drain.

 

Wk of 8/30/04:  ESI continues with waste concrete excavation, crushing & blending eng'd fill material.  They continue to work the levee fill from sta 26+00 to sta 32+00;  from sta 40+00 to sta 52+00;  & from sta 67+00 to sta 98+00.  They also continue to haul lev fill material from Simpson Lake Det area & from Simpson Material.  They also started installing sheet piling behind Mer Valley Plaza.  Loana continues to work on 3rd St gravity drain & on 60" RCP gravity drain from Pharoah west to Det area "D" behind the biz on the south of Marshall Rd from 7th St to 9th St.  Jacobsmeyer (JMC) continues to work on 54" gravity drain at BPlz.  Kelpe completed placing rip-rap, bedding & C-stone at 141 gravity drain Outlet Structure. 

 

Wk of 9/6/04:  ESI continues with waste concrete excavation, crushing & blending eng'd fill material.  They continue to work the levee fill from sta 26+00 to sta 32+00;  from sta 40+00 to sta 52+00;  & from sta 67+00 to sta 98+00.  They also continue to haul levee fill material from Simpson Lake Det area & from Simpson Material.  They completed installing the sheet piling behind Mer Valley Plaza.  Loana continues to work on the 60" RCP gravity drain from Pharoah west to Det area "D" behind the biz on the south of Marshall Rd from 7th St to 9th Sts.  JMC continues to work on 54" gravity drain at BPlz.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 10/18/04

Subj:  Item 4B (Phase 3)

 

Wk of 9/13/04:  ESI continues with waste concrete excavation, crushing & blending eng'd fill material.  They continue to work the levee fill from sta 26+00 to 32+00;  from sta 40+00 to sta 52+00;  from sta 67+00 to sta 98+00.  They also continue to haul levee fill material from Simpson Lake Det area & from Simpson Material.  Loana continues to work on 60"0 R.C.P. gravity drain from Pharoah west to Det Area "D".  JMC continues to work on 54"0 gravity drain outlet structure at BPlz.  They also continue to work on 3rd St Gatewell & Junction Box.  Sub-contractor Contract Dewatering started installing relief wells at Simpson Lake Det area.  MO American Water Co (MAWC) is installing water main east of Kena on Pyramid & under levee on Pyramid.  Bom Const Co has completed installing new MSD Lift sta on Marshall Rd & are waiting for MSD approval so they can remove the old existing lift sta on Kena. 

 

Wk of 9/20/04:  ESI continues with waste concrete excavation, crushing & blending eng'd fill material;  they continue to work the levee fill from sta 26+00 to 32+00;  from sta 40+00 to 52+00;  from sta 67+00 to 98+00;  they also continue to haul levee fill material from Simpson Lake Det area & from Simpson Material.  Loana continues to work on 60"0 R.C.P. gravity drain from Pharoah west to Det Area "D".  They also started installing various 0 {sic} R.C.P. gravity drain behind MerPlz. 

 

JMC continues to work on 3rd St Gatewell & Junction Box.  Contract Dewatering continues installing relief wells at Simpson Lake Det area.  MAWC continues installing water main east on Kena & Pyramid.  Bom Const Co is still awaiting final approval of a new lift sta by MSD so they can remove the old existing lift sta.

 

Wk of 9/27/04:  ESI continues with waste-concrete excavation & crushing ops.  They have completed the eng'g fill & blending ops.  They continue to work the levee fill from sta 26+00 to 32+00;  from sta 40+00 to 54+00;  from sta 57+00 to 98+00.  They have completed excavating & hauling from Simpson Lake Det area;  started to excavate & haul levee fill material for GG Det area. They continue to haul levee fill material from Simpson Const Material. 

 

Loana continues to work on the 60"0 R.C.P. gravity drain from Pharoah west to Det Area D & on the gravity drain behind MerPlz.  JMC continues to work on 3rd St Gatewell & Junction Box.  Contract Dewatering completed installing relief wells at Simpson Lake Det area & has started to install relief wells at Det Area C.  MAWC has completed installing water mains on Kena & Pyramid.  Bom is still awaiting final approval of the new lft sta by MSD. 

 

Wk of 10/4/04:  ESI continues with waste-concrete excavation & crushing ops.  They continue to work the lev fill from sta 26+00 to 32+00;  from sta 40+00 to 54+00;  from sta 57+00 to 98+00.  With fill material being hauled from GG Det area & from Simpson Const Material.  Loana continues to work on the 60"0 R.C.P. gravity drain from Pharoah west to Det Area D & has completed the gravity drain behind MerPlz.  JMC continues to work on 3rd St Gatewell.  Bom is still awaiting final approval of new lft sta by MSD.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 1/18/05

Subj:  Item 4B (Phase 3)

Wk of 12/6/04:  ESI continues to be held up on levee const due to wet conditions & rain.  Sub-JMC continues to work on flood wall on South side of MerVPlz.  Wk of 12/13/04:  ESI started hauling levee fill material from GG stockpile & Simpson Material.  Sub-JMC continues with flood wall const.  Wk of 12/20/04:  ((same as 12/13)).  Wk of 12/27/04:  ((same as 12/6 + for both))  weather permitting.  Wk of 1/3/05:  Progress on levee project halted due to wet conditions on site & rain.  >  DC.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 2/22/05

Subj:  Item 4B (Phase 3)

Wk of 1/3/05:  Due to rain on the previous wkend & early in the wk the Nat'l Weather Svc issued a river forecast of 24.00 ft (417.00) on 1/8.  This elevation would impact levee const.  On 1/5 a mtg was held with COE, COVP{sic} & ESI.  It was decided that ESI would fill the gap in the levee at Pharoah with levee fill material from the top of levee on the West side of Pharoah to a elevation of approx 27.00 ft (420.00) & placed stop logs in junction boxes at 3rd & 5th St outlet structures.  This would prevent the river from back flowing into the det ponds the City Emerg Op Crew closed the sluice gates at Fishpot Creek, GG Creek & Simpson Lake.  They also placed a stop log in junction box at Mer Sta Rd & BPlz to prevent storm water drainage from Hwy 141 from going into Fishpot det pond.  This work was all completed on 1/5.  The river crested at approx 23.00 (416.00) on 1/8.  A great job was done by ESI & the City Emerg Crew to complete this work in such a short time & protecting any damage to the existing levee const. 

 

Wk of 1/10/05:  Progress on levee project halted due to flooding & high water conditions.  The precautions taken last wk to protect lev const will remain in place until flooding conditions recede.

Wk of 1/17/05:  Progress on levee project is still on hold due to wet conditions.  ESI removed some of the stop logs at 3rd & 5th Sts to allow interior water from det ponds to drain to river.

Wk of 1/24/05:  Progress on levee is still on hold due to wet conditions.

Wk of 1/31/05:  Progress on levee is still on hold due to wet conditions.

Wk of 2/7/05:    Progress on levee is still on hold due to wet conditions.

Wk of 2/14/05:  Progress on levee is still on hold due to wet conditions.  >  DC: mw.

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 3/21/05

Subj:  Item 4B (Phase 3)

Wk of 2/21/05:  ESI started to haul stockpiled material from GG Det area & lay it out, disc & place on levee when material dried.  A progress mtg was held on 2/24 to discuss with COE, ESI & VP on pending issues that needed to be resolved before weather breaks & we have no delays in const.  Several pending issues need to be resolved, none will hold up const presently.

 

Wk of 2/28/05:  ESI continues to haul material from the stockpile at GG Det area.  Sub Contractor Kelpe started to work on storm water junction box on Mer Sta Rd south of Vance Rd, west of 141 & north of UPRR overpass. 

Wk of 3/7/05:  ESI continues to haul material from the stockpile at GG Det area.  They also started hauling material from Simpson Material.  Sub-contractor Kelpe continues to work on storm water junction box on Mer Sta Rd.  The relief well contractor is drilling wells along landside of levee. >  DC: mw 

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 4/18/05

Subj:  Item 4B (Phase 3)

Wk of 3/14/05:  ESI continues to haul material from Simpson Material.  They will continue to haul from the GG Det area stockpile when material dries out sufficiently to remove & haul.  Sub-Kelpe has completed work on storm water junction box on MerStaRd gates & operators are installed. Contractor CDS has completed drilling relief well & will perform relief well pump test in 2 wks. 

 

Wk of 3/21/05:  Due to the rain on the previous wkend & early in the wk progress on levee was put on hold.  Wk of 3/28/05:  Due to rain early in wk hauling of material was halted but with day {sic} weather later in wk, ESI was able to haul levee material from Simpson on Thurs & Fri.  ESI has completed removing the stockpiled levee material from GG Det area & will complete the finish work as weather permits.  Wk of 4/4/05:  ESI continues to haul material from Simpson Material & place on levee.  Sub-Kelpe has started to install the sluice gates in the gatewell structures at 3rd & 5th Sts.  Landscaping & tree planting has been completed in Simpson Lake Det area.  >  DC:mw

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 5/16/05

Subj:  Item 4B (Phase 3)

 

Wk of 4/11/05:  Due to rain this wk progress on levee was put on hold.  Wk of 4/18/05:  Due to inclemt weather this wk progress on levee was put on hold.  Wk of 4/25/05:  Due to inclemt weather this wk ESI is still unable to haul levee fill material from Simpson Material.  Sub-Kelpe has completed installing sluice gates at 3rd & 5th Sts & has started installing gate operators.  Wk of 5/2/05:  ESI has again started to haul levee fill material from Simpson Material (material is somewhat wet but contractor is able to spread & dry it out).  ESI has started to finish grade & seed the completed part of the AL Levee & the Fishpot Det area to reduce material erosion on levee slopes.  Sub-Kelpe continues to install gate operators at 3rd & 5th Sts.  Sub-Ford Asphalt Co, Inc started & completed the asphalt paving at BPlz pkg lot.  > DC:mw

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA - 5/16/05 {sic} ((should be 6/20/05))

Subj:  Item 4B (Phase 3)

 

Wk of 5/9/05:  ESI continues to haul levee fill material from Simpson Material.  Sub-Kelpe continues to install the gate operator {sic} at 3rd & 5th St outlet structures.  Wk of 5/16/05:  ESI continues to haul levee fill material from Simpson Material.  Sub-Kelpe continues to install gate operators {sic} at 3rd & 5th St outlet structures.  Wk of 5/23/05: ((same as 5/16 +)) City has obtained their ins & permits from BNSF so the storm water lines at MVP  {sic} ((surely MerVPlz)) can be started on.  Sub-Loana will start work on the line at the S side of MVP on approx 6/8/05.  When the work on this line is complete Sub-Kelpe will begin work on storm water line E of MVP along BNSF RR. 

 

Wk of 5/30/05:  ESI continues to haul levee fill material from Simpson Material & place on levee.  Sub-Loana has started to install the storm water piping into the 3rd St junction box (JB-2) at 3rd St.  They will start to work on the storm water line at the S side MVP on 6/8/05.  City has acquired verbal permission from Mr Paul Litteton {sic} of BNSF to use the trailer entrance at 101 Marshall for the haul road to the project.  EM has put in a request to BNSF for a perm crossing after project is complete.  Wk of 6/6/05:  ESI was only able to haul levee fill material from Simpson Material & place on levee for 3 days this wk due to rain on Wed & Thurs nights.  Sub-Leona has completed the storm water piping at 3rd St but was unable to start on the storm water line at MVP this wk due to rain.  >  DC:mw

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA -  9/19/05

Subj:  Item 4B (Phase 3)

 

Wk of 8/8/05:  ESI has completed placing levee fill material on the East side of Hwy 141 from Sta 33 + 80 to the StL Ave closure structure at Sta 104 + 80.  ESI continues to haul fill material to construct access ramp at Marshall Rd & Emerg Access ramp East of the BN RR ROW.  ESI has installed a layer of geotextile material on the levee road East of BN RR ROW tie in to the StL Ave tie in.  ESI has started to place thhe aggregate surfacing material on the above levee road.  Subcontractor Kelpe started to work on the new 54" diameter storm water pipe & sluice gate structure on the West side of Hwy 141.

Wk of 8/15/05:  ESI has completed placing the aggregate surfacing material on levee road.  The hauling & placing of fill material on levee was held up this wk to {sic} wet weather conditions.  Sub-Kelpe continued work on the storm water pipe West of Hwy 141 when weather permits. 

 

Wk of 8/22/05:  ESI started to seed & mulch the river side of levee from StL Ave Westward.  This op was halted on Wednesday due to rain;  all work by ESI has been stopped due to rain.  Sub-Kelpe continues to work on the storm water pipe West of Hwy 141 when weather permits.

Wk of 8/29/05:  ESI continues to seed & mulch the levee.  Sub-Kelpe has completed laying the new 54" diameter storm water pipe on the West side of Hwy 141.  They also have completed laying the 30" diameter storm water pipe to det area A (Arnold's Landing).  Sub-Jacobsmeyer started working on the 141 gatewell structure.  Sub-Ford Asphalt Paving completed the overlay paving of 9th St from StL Ave to Benton St.

Wk of 9/5/05:  ESI continues to seed & mulch the levee.  Sub-Jacobsmeyer continues to work on the 141 gatewell structure.  VP's Sub-Bom started working on the 300 foot 30" storm water line on the East side of MerVPlz from the existing manhole EX1-A {sic}  North to Marshall Rd.  DC:mw

_____________________________________________________________

To:  Mayor, VP Levee Cmsn, City Atty & BOA -  10/17/05

Subj:  Item 4B (Phase 3)

Wk of 9/12/05:  ESI continues to have {sic} & place levee fill material at various Stations on levee.  They continue to seed & mulch as weather permits.  Sub-Jacobsmeyer completed the concrete work on the 141 gatewell structure.  VP's Sub-Bom continues to work on the 300 foot 30" storm water line.  All work was halted this wk due to rain in the middle of the wk.

Wk of 9/19/05:  ESI continues to haul & place levee fill material at the 141 gatewell structure.  Also they are installing gate, stem & operator in gatewell structure.  Weather has halted all other work.

Wk of 9/26/05:  Work on levee was put on hold due to rain. 

Wk of 10/3/05:  ESI have {sic} completed hauling & placing levee fill material at Arnold's Landing.  They have also completed installing the sluice gate, stem & operator in the 141 gatewell structure.    DC:mw